OpenStudy (anonymous):

what is the restriction on the product of x^2+5x+4/x^2+3x-4*x^2+4x-5/x+3

OpenStudy (anonymous):

\[\frac{ x ^{2}+5x+4 }{x ^{2}+3x-4 }\times \frac{ x ^{2}+4x-5 }{x+3 }\] \[x ^{2}+3x-4=x ^{2}+4x-1x-4=x \left( x+4 \right)-1\left( x+4 \right)=\left( x+4 \right) (x-1)\] values of x for which denominator is 0 are -4,1,-3 Restrictions are \[x \neq -4,1,-3\]